# Test of the Equivalence Principle for Superconductors

### v2

@@ Abstract
We searched for violations of the weak equivalence principle using a cryogenic torsion balance with a pendulum comprised of superconducting niobium and copper. We constrain the Eötvös parameter with 95%-confidence to $\eta_{\text{Nb*-Cu}}~{}\leq~{}2.0\times 10^{-9}$ and $\eta_{\text{CP-ee}}\leq 6.9\times 10^{-3}$ for superconducting niobium and Cooper pairs, respectively.
@@
where $m_{e}$ and $m_{p}$ are respectively the mass of the electron and proton, $N_{e}$, $N_{n}$, $N_{v}$, and $N_{CP}$ are the number of electrons, nucleons, valence electrons, and electrons in Cooper pairs, respectively. The ratio $N_{CP}/N_{v}$ was estimated via Kittel and McEuen (2018):
$\frac{N_{CP}}{N_{v}}=\frac{m_{e}a^{3}}{2\mu_{0}e^{2}\lambda_{L}^{2}}$ (9)
where $a$ is the lattice constant, $\mu_{0}$ is the vacuum permeability, $e$ is the charge of the electron, and $\lambda_{L}$ is the London penetration depth.
[…]
@@
$\eta_{\text{CP-ee}}\leq 6.9\times 10^{-3}.$ (10)
Conclusion - We used a cryogenic torsion pendulum with superconducting-niobium and copper test bodies to constrain deviations from the WEP for superconducting niobium and Cooper pairs with 95%-confidence to $\eta_{\text{Nb*-Cu}}~{}\leq~{}2.0\times 10^{-9}$ and $\eta_{\text{CP-ee}}\leq 6.9\times 10^{-3}$, respectively. This improves upon the result reported by Tajmar et al. (2009) by more than two orders of magnitude. Our experiment places strong constraints on any enhanced gravitational effects of superconductors.
@@
Acknowledgements - Participation from the University of Washington, Seattle, was supported by funding from the NSF under Awards PHY-1607385, PHY-1607391, PHY-1912380, and PHY-1912514.

### v3

@@ Abstract
We searched for violations of the weak equivalence principle using a cryogenic torsion balance with a pendulum comprised of superconducting niobium and copper. We constrain the Eötvös parameter with 95%-confidence to $\eta_{\text{Nb*-Cu}}~{}\leq~{}2.0\times 10^{-9}$ and $\eta_{\text{CP-ee}}\leq 9.2\times 10^{-4}$ for superconducting niobium and Cooper pairs, respectively.
@@
where $m_{e}$ and $m_{p}$ are respectively the mass of the electron and proton, $N_{e}$, $N_{n}$, $N_{v}$, and $N_{CP}$ are the number of electrons, nucleons, valence electrons, and electrons in Cooper pairs, respectively. The ratio $N_{CP}/N_{v}$ was estimated with the two fluid model Sharma (2021):
$\frac{N_{CP}}{N_{v}}=1-\bigg{(}\frac{T}{T_{c}}\bigg{)}^{4}$ (9)
where $T$ is the temperature and $T_{c}$ is the critical temperature. We use a conservative upper limit of the temperature of the pendulum of 8 K which corresponds to a Cooper pair fraction of $N_{CP}/N_{v}=0.38$.
@@
$\eta_{\text{CP-ee}}\leq 9.2\times 10^{-4}.$ (10)
Conclusion - We used a cryogenic torsion pendulum with superconducting-niobium and copper test bodies to constrain deviations from the WEP for superconducting niobium and Cooper pairs with 95%-confidence to $\eta_{\text{Nb*-Cu}}~{}\leq~{}2.0\times 10^{-9}$ and $\eta_{\text{CP-ee}}\leq 9.2\times 10^{-4}$, respectively. This improves upon the result reported by Tajmar et al. (2009) by more than two orders of magnitude. Our experiment places strong constraints on any enhanced gravitational effects of superconductors.
@@
Acknowledgements - Special thanks to Andreas Schilling for help with superconductor physics. Participation from the University of Washington, Seattle, was supported by funding from the NSF under Awards PHY-1607385, PHY-1607391, PHY-1912380, and PHY-1912514.